# GUI-World: A Dataset for GUI-oriented Multimodal LLM-based Agents

### v1

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have been used as agents to control keyboard and mouse inputs by directly perceiving the Graphical User Interface (GUI) and generating corresponding code. However, current agents primarily exhibit excellent understanding capabilities in static environments and are predominantly applied in relatively simple domains, such as Web or mobile interfaces. We argue that a robust GUI agent should be capable of perceiving temporal information on the GUI, including dynamic Web content and multi-step tasks. Additionally, it should possess a comprehensive understanding of various GUI scenarios, including desktop software and multi-window interactions. To this end, this paper introduces a new dataset, termed GUI-World, which features meticulously crafted Human-MLLM annotations, extensively covering six GUI scenarios and eight types of GUI-oriented questions in three formats. We evaluate the capabilities of current state-of-the-art MLLMs, including ImageLLMs and VideoLLMs, in understanding various types of GUI content, especially dynamic and sequential content. Our findings reveal that ImageLLMs struggle with dynamic GUI content without manually annotated keyframes or operation history. On the other hand, VideoLLMs fall short in all GUI-oriented tasks given the sparse GUI video dataset. Based on GUI-World, we take the initial step of leveraging a fine-tuned VideoLLM as a GUI agent, demonstrating an improved understanding of various GUI tasks. However, due to the limitations in the performance of base LLMs, we conclude that using VideoLLMs as GUI agents remains a significant challenge. We believe our work provides valuable insights for future research in dynamic GUI content understanding. The code and dataset are publicly available at our project homepage: https://gui-world.github.io/.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs), such as GPT-4V(ision) [1] and LLaVA [2], have significantly contributed to the development of the visual-text domain [3]. These models bring forth innovative solutions and paradigms for traditional visual tasks, including visual reasoning [4], medical image interpretation [5, 6], and applications in embodied agents [7]. One particularly promising area is Graphical User Interface (GUI) understanding, which holds significant potential for real-world applications, such as webpage comprehension [8, 9] and navigation by GUI agents [10, 11, 12]. The key challenges of GUI understanding are twofold: effective GUI agents are expected to (1) possess a deep understanding of GUI elements, including webpage icons, text identified through Optical Character Recognition (OCR), and page layouts, and (2) exhibit an exceptional ability to follow instructions within GUI contexts, such as conducting searches through search engines.
@@
Despite significant progress, as illustrated in Table 1, existing works suffer from the following limitations: (1) Most studies predominantly focus on the static features of GUI scenarios, neglecting the need for MLLMs to effectively process sequential information and dynamic operations. For instance, an agent’s task performance can be disrupted by unexpected elements such as pop-up advertisements, underscoring a gap in handling dynamic sequential tasks. (2) Current research is typically restricted to Web-based environments, which limits the models’ generalization and robustness. For instance, GUI agents may need to operate across diverse platforms such as Windows, macOS, Linux, iOS, Android, and XR environments. Additionally, operations may sometimes involve multiple windows. Therefore, expanding the scope of research to encompass these varied environments will enhance the adaptability and effectiveness of GUI agents.
@@
To mitigate these gaps, this paper introduces GUI-World, a comprehensive dataset containing over 12,000 GUI videos, specifically designed to evaluate and enhance the capabilities of GUI agents. This dataset encompasses a wide range of GUI scenarios, including popular websites, desktop and mobile applications across various operating systems, multi-window interactions, as well as XR environments. The data collection process involves sourcing GUI videos from screen recordings and instructional videos on YouTube. Subsequently, we utilize an Human-MLLM collaborative approach to generate a diverse set of questions and instructions and finally construct GUI-World.
@@
| Rico [13] | 72,219 | Low | ✔ | ✔ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | UI Code/Layout Generation |
| MetaGUI [14] | 1,125 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | Mobile Navigation |
| UGIF [15] | 523 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | UI Grounded Instruction Following |
| AITW [16] | 715,142 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | GUI Understanding |
| Ferret-UI [17] | 123,702 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | UI Grounding & Understanding |
| MiniWoB++ [18] | 100 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | ✘ | Web Navigation |
| WebArena [19] | 812 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| Mind2Web [20] | 2,350 | Both | ✔ | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| OmniAct [21] | 9,802 | Low | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✘ | Code Generation |
| MMINA [22] | 1,050 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✔ | ✘ | Web Navigation |
| AgentStudio [23] | 304 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
| OSWorld [24] | 369 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
@@
Likewise, we also establish a comprehensive benchmark for GUI understanding, which encompasses seven mainstream MLLMs, three keyframe selection strategies, six GUI scenarios, and a diverse array of queries in multiple-choice, free-form, and conversational formats, aiming to provide a thorough evaluation of the MLLMs’ GUI-oriented capabilities. As shown in Figure 2, the assessment results indicate that most MLLMs struggle with GUI-World, highlighting their limited dynamic understanding of graphical interfaces and underscoring the need for further enhancement.
Leveraging this dataset, we take the first step of fine-tuning a Video GUI Agent proficient in dynamic and sequential GUI tasks, which results in significant improvements in the general capabilities of GUI agents, thereby demonstrating the utility and effectiveness of GUI-World. Additionally, we delve into discussing various factors critical to GUI understanding, including the integration of textual information, the number of keyframes, and image resolutions.
Overall, the key contributions of this paper are three-fold:
A New Dataset. We propose GUI-World, a comprehensive GUI dataset comprising over 12,000 videos specifically designed to assess and improve the GUI understanding capabilities of MLLMs, spanning a range of categories and scenarios, including desktop, mobile, and extended reality (XR), and representing the first GUI-oriented instruction-tuning dataset in the video domain.
A Novel Model. Based on GUI-World, we propose GUI-Vid, a GUI-oriented VideoLLM with enhanced capabilities to handle various and complex GUI tasks. GUI-Vid shows a significant improvement on the benchmark and achieves results comparable to the top-performing models.
Comprehensive Experiments and Valuable Insights. Our experiments indicate that most existing MLLMs continue to face challenges with GUI-oriented tasks, particularly in sequential and dynamic GUI content. Empirical findings suggest that improvements in vision perception, along with an increase in the number of keyframes and higher resolution, can boost performance in GUI-oriented tasks, thereby paving the way for the future of GUI agents.
2 GUI-World: A Comprehensive Dataset for GUI Understanding
@@ 2.1 Overview
We introduce GUI-World, a comprehensive dataset covering six GUI scenarios including video, human-annotated keyframes, as well as detailed captions and diverse types of QA produced by our data curation framework, aiming at benchmarking and enhancing the general GUI-oriented capabilities. These GUI scenarios encompass desktop operating systems (e.g., macOS, Windows) and mobile platforms (e.g., Android and iOS), websites, software, and even extended-range technologies (XR) (e.g., GUI in Apple Vision Pro [25]). Discussion for each scenario is in subsection A.1.
As illustrated in Figure 3, the development of GUI-World is structured around a two-stage process. Details regarding video and query statistics are provided in Table 2, which includes distributions of the number of keyframes, video lengths, and the lengths of queries and their corresponding golden answers, as displayed in Figure 4. Refer to Figure 5 and Appendix F for case study.
@@ 2.2 GUI Video Collection and Keyframe Annotation Process
Despite the high fidelity of these manually recorded videos, we encounter several challenges: (1) Student workers often require substantial time to acquaint themselves with professional software (e.g., MATLAB, Adobe After Effects (Ae)), which can hinder the progress of data collection. (2) The videos may lack comprehensiveness, typically capturing only commonly used operations and overlooking rarer functions crucial for dataset completeness. To address these issues, we also source videos from social media platforms that host a diverse array of GUI-related content. Specifically, we download tutorial videos from YouTube—given its prevalence as a video-sharing platform—because they richly detail various GUI operations. These videos are then segmented into shorter clips, each representing a distinct sequence of operations.
The subsequent step involves annotating these video clips with keyframes and textual descriptions of each keyframe using custom-designed annotation software. Although several algorithms exist for keyframe extraction [26, 27, 28, 29], they typically underperform with GUI videos where changes between frames might be minimal (e.g., a slight movement in the mouse cursor). To ensure high-quality datasets, we therefore perform manual extraction of these keyframes. Each keyframe is meticulously annotated to include details such as the operation performed, the purpose between two keyframes, the software or website used, mouse actions (e.g., scroll, click), and keyboard inputs (e.g., copy (Ctrl + C), paste (Ctrl + V), specific input). We detail our annotation process in subsection A.3.
@@ 2.3 GUI Tasks Generation from Human-MLLM Collaboration
Drawing insights from prior research [30, 31, 32, 33, 34], we develop a Human-MLLM collaboration pipeline to annotate captions and diverse types of QA specifically tailored for GUI comprehension. The process involves inputting an instructional prompt, a comprehensive description, key information (e.g., system or application), and a sequence of human-annotated keyframes into GPT-4V. As depicted in Table 3, GUI-World features an array of question types, as detailed in follows:
Detailed and Summarized Captioning: This task challenges basic GUI knowledge and multimodal perception, also addressing the deficiency of detailed GUI content in video-caption pairs. Initially, GPT-4V generates two distinct descriptions for each video: one concentrating on fine-grained details and the other on the overall image sequences. Furthermore, GPT-4V provides a succinct summary, highlighting core operations and overarching objectives in the video.

### v2

# GUI-World: A Video Benchmark and Dataset for Multimodal GUI-oriented Understanding

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have been used as agents to control keyboard and mouse inputs by directly perceiving the Graphical User Interface (GUI) and generating corresponding commands. However, current agents primarily demonstrate strong understanding capabilities in static environments and are mainly applied to relatively simple domains, such as Web or mobile interfaces. We argue that a robust GUI agent should be capable of perceiving temporal information on the GUI, including dynamic Web content and multi-step tasks. Additionally, it should possess a comprehensive understanding of various GUI scenarios, including desktop software and multi-window interactions. To this end, this paper introduces a new dataset, termed GUI-World, which features meticulously crafted Human-MLLM annotations, extensively covering six GUI scenarios and eight types of GUI-oriented questions in three formats. We evaluate the capabilities of current state-of-the-art MLLMs, including Image LLMs and Video LLMs, in understanding various types of GUI content, especially dynamic and sequential content. Our findings reveal that current models struggle with dynamic GUI content without manually annotated keyframes or operation history. On the other hand, Video LLMs fall short in all GUI-oriented tasks given the sparse GUI video dataset. Therefore, we take the initial step of leveraging a fine-tuned Video LLM, GUI-Vid, as a GUI-oriented assistant, demonstrating an improved understanding of various GUI tasks. However, due to the limitations in the performance of base LLMs, we conclude that using video LLMs as GUI agents remains a significant challenge. We believe our work provides valuable insights for future research in dynamic GUI content understanding. All the dataset and code are publicly available at: https://gui-world.github.io.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs), such as GPT-4V(ision) (OpenAI, 2023) and LLaVA (Liu et al., 2023b), have significantly contributed to the development of both vision and language domains (Yin et al., 2024). These models bring forth innovative solutions and paradigms for traditional visual tasks, including visual reasoning (Yang et al., 2023b), medical image interpretation (Li et al., 2024b), and applications in embodied agents (Huang et al., 2023). One particularly promising area is Graphical User Interface (GUI) understanding, which holds significant potential for real-world applications, such as webpage comprehension (Hong et al., 2024, Lai et al., 2024) and navigation by GUI agents (Yang et al., 2023a, Niu et al., 2024, Wang et al., 2024). The key challenges of GUI understanding are twofold: effective GUI agents are expected to (1) possess a deep understanding of GUI elements, including webpage icons, text identified through Optical Character Recognition (OCR), and page layouts, and (2) exhibit an exceptional ability to follow instructions within GUI contexts, such as conducting searches through search engines.
@@
Despite significant progress, as illustrated in Table 1, prior studies on GUI-related datasets and benchmarks suffer the following limitations: (1) Inability to Handle Dynamic Environments. Most studies primarily emphasize the static features of GUI scenarios, often overlooking the importance of enabling MLLMs to effectively handle dynamic information and sequential operations. For instance, an agent’s task performance can be disrupted by unexpected elements such as pop-up advertisements, underscoring a gap in handling dynamic sequential tasks. (2) Limited Scenarios. Current research is typically restricted to Web-based or mobile environments, which limits to assess their generalization ability and robustness across scenarios. For instance, GUI-oriented models may need to operate across diverse platforms such as Windows, macOS, Linux, iOS, Android, and eXtended Reality (XR) environments. Operations may sometimes involve multiple windows. Therefore, expanding research to encompass these varied environments will enhance the adaptability and effectiveness.
@@
To mitigate these gaps, this paper introduces GUI-World, a comprehensive dataset containing 12,379 GUI videos, specifically designed to evaluate and enhance the capabilities of GUI agents. This dataset encompasses a wide range of GUI scenarios, including popular websites, desktop and mobile applications across various operating systems, multi-window interactions, as well as XR environments. The data collection process involves sourcing GUI videos from screen recordings and instructional videos on YouTube. Subsequently, we utilize a Human-MLLM collaborative approach to generate a diverse set of captions, complex queries, and multi-round conversation, ultimately constructing GUI-World.
Likewise, we also establish a comprehensive benchmark for GUI understanding, which encompasses nine mainstream MLLMs (e.g., GPT-4o (OpenAI, 2024) and Gemini-1.5-Pro (Team et al., 2023)), five keyframe selection strategies (e.g., UVD (Zhang et al., 2024d)), and six GUI scenarios, aiming to provide a thorough evaluation of the GUI-oriented understanding capabilities of MLLMs. As shown in Section 2, the assessment results indicate that most MLLMs struggle with GUI-World, highlighting their limited dynamic understanding of graphical interfaces and underscoring the need for further enhancement.
Using this dataset, we take the first step to fine-tune a GUI-oriented Video LLM that excels at handling dynamic and sequential GUI tasks, leading to substantial enhancements in the general capabilities and showcasing the utility and effectiveness of GUI-World. Additionally, we delve into discussing various factors critical to GUI understanding, including the integration of textual information, the number of keyframes, image resolutions, and vision perception, providing a pioneering and comprehensive study of the GUI domain.
Overall, the key contributions of this paper are threefold:
A New Dataset. We propose GUI-World, a comprehensive GUI dataset comprising 12,379 videos specifically designed to assess and improve GUI-oriented capabilities of MLLMs, spanning a range of categories and scenarios, including desktop, mobile, and XR environments. It stands as the first GUI-oriented instruction-tuning dataset in video domain.
Comprehensive Experiments and Valuable Insights. Our experiments indicate that most existing MLLMs continue to face challenges with GUI-oriented tasks, particularly in sequential and dynamic GUI content. Empirical results suggest that enhancing vision perception, such as more keyframes and higher resolution, can lead to substantial performance improvements in GUI tasks.
A Explorative GUI-oriented Video LLM. Based on GUI-World, we propose GUI-Vid, a GUI-oriented video LLM with enhanced capabilities to handle various and complex GUI tasks. GUI-Vid shows a significant improvement on the benchmark and achieves results comparable to the top-performing models, thereby paving ways for the future of GUI models.
@@
| Rico (Deka et al., 2017) | 72,219 | Low | ✔ | ✔ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | UI Code/Layout Generation |
| MiniWoB++ (Liu et al., 2018) | 100 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | ✘ | Web Navigation |
| Screen2Words (Wang et al., 2021) | 22,417 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | UI Summarization |
| MetaGUI (Sun et al., 2022) | 1,125 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | Mobile Navigation |
| UGIF (Venkatesh et al., 2022) | 523 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | Instruction Following |
| AITW (Rawles et al., 2023) | 715,142 | High | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | GUI Understanding |
| Ferret-UI (You et al., 2024) | 123,702 | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | UI Grounding & Understanding |
| Spotlight (Li & Li, 2022) | 2.5M | Low | ✔ | ✘ | ✘ | ✔ | ✘ | ✘ | ✘ | ✘ | ✘ | GUI Understanding |
| WebArena (Zhou et al., 2023) | 812 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| Mind2Web (Deng et al., 2024) | 2,350 | Both | ✔ | ✔ | ✔ | ✘ | ✘ | ✘ | ✔ | ✘ | ✘ | Web Navigation |
| OmniAct (Kapoor et al., 2024) | 9,802 | Low | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✘ | Code Generation |
| GUICourse (Chen et al., 2024c) | 10.7M | Both | ✔ | ✘ | ✔ | ✔ | ✘ | ✘ | ✔ | ✔ | ✘ | GUI Understanding |
| MMINA (Zhang et al., 2024e) | 1,050 | Low | ✔ | ✘ | ✔ | ✘ | ✘ | ✘ | ✔ | ✔ | ✘ | Web Navigation |
| AgentStudio (Zheng et al., 2024b) | 304 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
| OSWorld (Xie et al., 2024) | 369 | High | ✔ | ✘ | ✔ | ✘ | ✔ | ✘ | ✔ | ✔ | ✘ | General Control |
@@
2 GUI-World: A Dataset for GUI Understanding
@@ 2.1 Overview
We introduce GUI-World, a comprehensive dataset covering six GUI scenarios including video, human-annotated keyframes, as well as detailed captions and diverse types of QA produced by our data curation framework, aiming at benchmarking and enhancing the general GUI-oriented capabilities. These GUI scenarios encompass desktop operating systems (e.g., macOS, Windows) and mobile platforms (e.g., Android and iOS), websites, software, and even XR (e.g., GUI in Apple Vision Pro (Apple, 2024)). We divide the dataset into a train-test split, each containing 10,702 and 1,677 samples. Discussion for each scenario is referred to Appendix A.1.
As illustrated in Figure 3, the development of GUI-World follows a two-stage process. Further details about video and query statistics are outlined in Table 2, including distributions of keyframe counts, video durations, query lengths, and their corresponding golden answers. Figures 4 and 5 further illustrate these data distributions.
@@ 2.2 GUI Video Collection and Keyframe Annotation Process
Despite the high fidelity of these manually recorded videos, we encounter several challenges: (1) Student workers often require substantial time to acquaint themselves with professional software (e.g., MATLAB, Adobe After Effects (Ae)), which can hinder the progress of data collection. (2) The videos may lack comprehensiveness, typically capturing only commonly used operations and overlooking rarer functions crucial for dataset completeness. To address these issues, we also source videos from social media platforms that host a diverse array of GUI videos. Specifically, we download tutorial videos from YouTube—given its prevalence as a video-sharing platform—because they richly detail various GUI operations. These videos are then segmented into shorter clips, each representing a distinct sequence of operations.
The subsequent step involves annotating these video clips with keyframes and textual descriptions of each keyframe using custom-designed annotation software. Although several algorithms exist for keyframe extraction (Zhu et al., 2016, Yan et al., 2018, Mahasseni et al., 2017, OpenCV, ), they typically underperform with GUI videos where changes between frames might be minimal (e.g., a slight movement in the mouse cursor). To ensure high-quality datasets, we therefore perform manual extraction of these keyframes. Each keyframe is meticulously annotated to include details such as the operation performed, the purpose between two keyframes, the software or website used, mouse actions (e.g., scroll, click), and keyboard inputs (e.g., copy (Ctrl + C), paste (Ctrl + V), specific input). We detail our annotation process in Appendix A.3.
@@ 2.3 GUI Tasks Generation from Human-MLLM Collaboration
Drawing insights from prior research (Dekoninck et al., 2024), we develop a Human-MLLM collaboration pipeline to annotate captions and diverse types of QA specifically tailored for GUI comprehension. The process involves inputting an instructional prompt, a comprehensive description, key information (e.g., system or application), and a sequence of human-annotated keyframes into GPT-4V. As depicted in Table 11, GUI-World features various question types, detailed as follows:
Detailed and Summarized Captioning: This task challenges basic GUI knowledge and multimodal perception, also addressing the deficiency of detailed GUI content in video-caption pairs. Initially, GPT-4V generates two distinct descriptions for each video: one concentrating on fine-grained details and the other on overall information. Furthermore, GPT-4V provides a succinct summary, highlighting core operations and overarching objectives in the video.